OpenStudy (anonymous):

What should be the axis of symmetry of the quadratic function f(x)=5x^2-6?

OpenStudy (raden):

use the formula : x= -b/2a

OpenStudy (raden):

here given b=0 and a=5 (just remember the general form quadratic fuction y=ax^2+bx+c)

OpenStudy (anonymous):

the choices are a. x=5 b. x=-5 c. x=-6 d. x=6

OpenStudy (raden):

hmmm.. x=-b/2a = -0/(2*5) = 0 are sure u typed correct options/the equation above ?

OpenStudy (anonymous):

yeah.......

OpenStudy (anonymous):

so this means that this is another typographical error?

OpenStudy (anonymous):

yes there is an error.

OpenStudy (anonymous):

ok thanks a lot